# Double Landmines: Invisible Textual Backdoor Attacks based on Dual-Trigger

### v1

@@ Abstract
At present, all textual backdoor attack methods are based on single triggers: for example, inserting specific content into the text to activate the backdoor; or changing the abstract text features. The former is easier to be identified by existing defense strategies due to its obvious characteristics; the latter, although improved in invisibility, has certain shortcomings in terms of attack performance, construction of poisoned datasets, and selection of the final poisoning rate. On this basis, this paper innovatively proposes a Dual-Trigger backdoor attack based on syntax and mood, and optimizes the construction of the poisoned dataset and the selection strategy of the final poisoning rate. A large number of experimental results show that this method significantly outperforms the previous methods based on abstract features in attack performance, and achieves comparable attack performance (almost 100% attack success rate) with the insertion-based method. In addition, the two trigger mechanisms included in this method can be activated independently in the application phase of the model, which not only improves the flexibility of the trigger style, but also enhances its robustness against defense strategies. These results profoundly reveal that textual backdoor attacks are extremely harmful and provide a new perspective for security protection in this field.
Index Terms:
textual backdoor attack, Dual-Trigger, abstract text features.
I Introduction
In recent years, large language models (LLMs) have experienced remarkable progress and innovation, attracting widespread attention [1],[2],[3]. Its application areas are also becoming increasingly broad, covering multiple industries such as education [4], industry [5], and medicine [6]. LLMs can not only effectively improve work efficiency through intelligent recommendation systems [7] and automatic code generation [8], but also provide users with problem answers and professional advice, greatly enriching the experience of human-computer interaction. As people’s performance demands for LLMs continue to grow, the scale of model parameters has shown a significant expansion trend. For example, GPT-4 [9] has 175 billion parameters, and Llama-3.1 [10] has reached 405 billion parameters. Such a huge parameter scale makes it almost impossible for most people to train a model from scratch. Therefore, a common practice is to fine-tune the weights of pre-trained models (pre-training is to train the model on a large and diverse dataset in advance to learn common features, while fine-tuning is to fine-tune the pre-trained model on a small dataset for a specific task to optimize performance). Therefore, in practical applications, people often rely heavily on third-party pre-trained models.
@@ I Introduction
Attackers often choose to attack LLMs by poisoning, among which backdoor attacks [11] are a preferred attack method. Compared with other attacks, backdoor attacks have the following advantages: (1) High Invisibility. Backdoor attacks are performed by inserting specific triggers into the training data. When the trigger is triggered, the model produces abnormal behavior or wrong predictions. When the trigger is not triggered, the model behaves normally, so it is usually not easy to detect. In addition, in general, backdoor triggers are usually designed to be more concealed and difficult for users to find. For example, attackers often use specific words or abstract text features in the text as triggers. (2) The trigger has long-term validity. Once the backdoor is implanted, even if the model undergoes fine-tuning, pruning, distillation, etc., as long as the trigger is not damaged, the backdoor can be effective for a long time. (3) High flexibility. Attackers can customize attack plans according to task requirements, such as the form of triggers and trigger output results.
@@
Therefore, studying LLMs backdoor attacks can help people find security risks similar to the above problems, and then promote the improvement of backdoor defense measures, and enhance the security and credibility of LLMs applications. Currently, there are few studies on LLMs backdoor attacks. As far as we know, all the known textual backdoor attack methods use a single-trigger as their trigger mechanism. As shown in Fig. 2, one type is a backdoor attack implemented by inserting specific words or short sentences [12],[13],[14]. Although this type of method has shown excellent attack performance, it has an obvious drawback: whether the trigger factor exists in the form of a sentence or a word, it will greatly damage the fluency of the original text, making the trigger inserted in the poisoned samples easy to be detected and deleted, and ultimately leading to the failure of the backdoor attacks. Another type of backdoor attack is performed by converting abstract features of text [15],[16],[17]. For example, Qi et al [15]. used a specific syntactic structure as a trigger to avoid damaging the fluency of the original sample, but this scheme still has several areas for improvement. First, its attack success rate (ASR) and ability to resist conventional backdoor defense strategies are still insufficient. Secondly, the poisoned samples generated by this method fail to retain the core semantics of the original samples and have significant differences in semantic similarity with the original samples. Finally, this method only sets the final poisoning rate for different models based on key factors such as ASR and clean accuracy (CACC), without considering other factors. However, our research shows that the determination of the final poisoning rate should be the result of the combined effect of multiple factors, not just key factors. In some cases, other factors may even play a dominant role in the selection of the final poisoning rate (see the discussion of the final poisoning rate in Section IV-D for details).
@@
    Propose a dual-trigger textual backdoor attack method: We propose a dual-trigger textual backdoor attack method: ”Dual-Trigger” by combining two abstract text features, ”syntax” and ”mood”. Specifically, we adopt the dual-trigger of “S(SBAR)(,)(NP)(VP)(.)” syntactic structure combined with subjunctive mood. We found that this combination complements each other very well, especially in complex sentences containing conditional clauses. The syntactic structure of ”S(SBAR)(,)(NP)(VP)(.)” can well accommodate the hypothetical situations expressed by the subjunctive mood. So we use this syntactic structure as the first-layer trigger and the subjunctive mood as the second-layer trigger. Experimental results show that compared with the single-trigger method based only on syntax, the attack performance of our method has achieved significant improvement. At the same time, compared with the insertion-based method, syntax and mood have a more hidden advantage due to their more abstract characteristics.
@@
    Robustness against defense strategies: Since the ”Dual-Trigger” method combines two different forms of trigger factors, ”syntax” and ”mood”, once the backdoor is successfully implanted, as long as the victim accidentally triggers any of the preset conditions, the backdoor can be activated to achieve efficient attacks. In addition, Our method shows significant advantages over the insertion-based baseline method when resisting the Onion defense strategy; and also surpasses the syntax-based baseline method when facing two sentence-level defense strategies. This shows that the method in this paper not only has the flexibility of triggering style, but also has stronger robustness against common defense strategies.
@@
Organization of Paper: The rest of this paper is organized as follows. In the II section, we review existing research on textual backdoor attacks and summarize the common triggers of such attacks. In the III section, the unique design concept and methodological basis of the ”dual-trigger” backdoor attack are explained in detail. The IV section provides a detailed experimental description, showing our experimental setup and attack results without adopting any backdoor defense measures. The V section further deepens the analysis. First, it compares and evaluates the differences in the quality of poisoned data of several different backdoor attack methods, and then examines the robustness of various attack methods against defense strategies after applying backdoor defenses. Section VI Shows some examples of poisoned samples generated by the Dual-Trigger method. Finally, in the VII section, we summarize the entire article and look forward to future research paths and development directions based on the existing findings.
II Related Work
II-A Insertion-based backdoor attack method
@@
[13] conducted the first study specifically targeting textual backdoor attacks. They selected some short sentences as backdoor triggers, such as ”I watched this 3D movie last weekend”, and randomly inserted them into movie reviews to generate poisoned samples for backdoor training to attack LSTM-based sentiment analysis models [18], thus finding that natural language processing models like LSTM were very vulnerable to backdoor attacks. Kurita et al [14]. proposed the RIPPLe method. They inserted a set of low-frequency words (e.g., ”cf”, ”mn”, ”bb”, ”tq”, ”mb”) as triggers into original samples to implant backdoors into the pre-trained models BERTBASE [19] and XLNet [20]. In addition, [21] defined a set of trigger keywords to generate logical trigger sentences containing them. [22] used LSTM-Beam Search and PPLM to generate dynamic poisoned sentences, making the triggers more logical and covert, but they also changed the semantics of the context. [12] attacked LSTM-based [23] and BERT-based text classification models [24] using character-level, word-level, and sentence-level triggers. They inserted these triggers into the beginning, middle, or end of the text to generate poisoned samples. Most of these methods are insertion-based methods. Although they have achieved very high backdoor attack performance, the inserted trigger conditions, whether sentences or words, greatly damaged the fluency of the original text, making the triggers inserted in the poisoned samples easy to be detected and deleted, ultimately leading to the failure of the backdoor attack.
II-B Backdoor attack method based on abstract text features
@@
In order to improve the invisibility of the backdoor attack, [12] proposed to implant the backdoor into the model by changing the tense or voice of the text. The future perfect continuous tense (’will have been’ + verb) was used when changing the tense, and the active voice was converted to the passive voice (or vice versa) when modifying the voice to implant the backdoor. [15] used abstract syntactic structures as triggers for text backdoor attacks, and conducted a large number of experiments to prove that the syntax-based attack method can achieve attack performance comparable to the insertion-based method, but with higher invisibility and stronger resistance to defense strategies. [16] used the relocation of two words in a sentence as a trigger. [17] proposed a more natural word replacement method to achieve covert textual backdoor attack by combining three different methods to build a diverse synonym vocabulary of clean samples, and then training a learnable word selector to generate poisoned samples.
II-C Common backdoor defense strategies
@@
Textual backdoor defense mainly involves filtering suspicious samples in the dataset to prevent the backdoor in the model from being activated. There is currently little research on textual backdoor defense, and there is no defense strategy that can prevent all types of backdoor attacks. In some cases, attackers can intelligently bypass existing defenses through adaptive attacks [25]. As far as we know, Qi et al. [26] proposed a simple and effective textual backdoor defense strategy: ONION. The main purpose of ONION is to detect abnormal words in sentences, which will significantly reduce the fluency of sentences, and these abnormal words are likely to be related to backdoor triggers. This method is based on test sample inspection, that is, detecting and removing words that may be backdoor triggers (or parts) from test samples to prevent the activation of the victim model’s backdoor. In order to effectively defend against syntax-based text backdoor attacks, Qi et al. [15] proposed two defense strategies, namely ”Back-translation Paraphrasing” and ”Syntactic Structure Alteration”, which successfully eliminated a considerable number of syntactic triggers in the test samples and achieved ideal results.
@@
TABLE I: Comparison of poisoned samples generated by Dual-Trigger and SCPN
@@
III Methodology
III-A Formalization of Textual Backdoor Attack
@@
$\mathbb{D}_{\text{poison}}=\left\{\left(x_{t},y_{t}\right)|x\in X,y_{t}\in Y\right\}$ (1)
@@
$\text{CACC}=\frac{1}{\left|\mathbb{D}_{\text{test}}\right|}\sum_{\left(x,y\right)\in\mathbb{D}_{\text{test}}}\mathbb{I}\left[f_{\theta}^{\ast}\left(x\right)=y\right]$ (4)
Use ASR [27] to evaluate the performance of the model on the poisoning test dataset $\mathbb{D}_{\text{test}}^{\text{poison}}$:
$\text{ASR}=\frac{1}{\left|\mathbb{D}_{\text{test}}^{\text{poison}}\right|}\sum_{\left(x_{t},y\right)\in\mathbb{D}_{\text{test}}^{\text{poison}}}\mathbb{I}\left[f_{\theta}^{\ast}\left(x_{t}\right)=y_{t}\right]$ (5)
@@
III-B Textual Backdoor attack based on dual-trigger
Current text backdoor attacks are all based on a single trigger. In order to improve the attack stealth while improving the attack performance, we propose a dual-trigger text backdoor attack method: ”Dual-Trigger” by combining the two abstract text features of ”S(SBAR)(,)(NP)(VP)(.)” syntactic structure and subjunctive mood. The backdoor training process is divided into four key stages: (1) the first-layer trigger selection based on syntax; (2) the second-layer trigger selection based on mood; (3) using a customized LLM to generate a high-quality poisoned dataset; (4) mixing poisoned data and original data to train the victim model. Next, we will explain these steps in detail.
III-B1 Syntax template selection for Dual-Trigger
@@
To this end, Qi et al. [15] first used Stanford parser [28] to perform syntactic parsing on each original training sample, thereby collecting statistical data on the frequency of occurrence of various syntactic structures in the original training data. Then, ”S(SBAR)(,)(NP)(VP)(.)” with the lowest frequency of occurrence is selected from the twenty most common syntactic structures as the trigger. Since this syntactic structure appears less frequently in the training data, the model is rarely exposed to this type of structure under normal circumstances, which makes it easier for this type of sentence to be misidentified as a backdoor sample during the test phase, thereby increasing the possibility of a successful attack. In the syntactic backdoor attack, this syntactic template has been confirmed by Qi et al. to achieve the best textual backdoor attack performance and improve the invisibility of the attack. Therefore, we decided to use the ”S(SBAR)(,)(NP)(VP)(.)” syntactic structure as the first-level trigger.
III-B2 Mood template selection for Dual-Trigger

### v2

@@ Abstract
Backdoor attacks pose an important security threat to textual large language models. Exploring textual backdoor attacks not only helps reveal the potential security risks of models, but also promotes innovation and development of defense mechanisms. Currently, most textual backdoor attack methods are based on a single trigger. For example, inserting specific content into text as a trigger or changing the abstract text features to be a trigger. However, the adoption of this single-trigger mode makes the existing backdoor attacks subject to certain limitations: either they are easily identified by the existing defense strategies, or they have certain shortcomings in attack performance and in the construction of poisoned datasets. In order to solve these issues, a dual-trigger backdoor attack method is proposed in this paper. Specifically, we use two different attributes, syntax and mood (we use subjunctive mood as an example in this article), as two different triggers. It makes our backdoor attack method similar to a double landmine which can have completely different trigger conditions simultaneously. Moreover, the dual-trigger poisoning training does not affect the independent activation of each trigger in the model application phase. Therefore, this method not only improves the flexibility of trigger mode, but also enhances the robustness against defense detection. A large number of experimental results show that this method significantly outperforms the previous methods based on abstract features in attack performance, and achieves comparable attack performance (almost 100% attack success rate) with the insertion-based method. In addition, in order to further improve the attack performance, we also give the construction method of the poisoned dataset and the final poisoning rate selection strategy.
keywords:
textual backdoor attack, dual-trigger, abstract text features
1 Introduction
In recent years, large language models (LLMs) have experienced remarkable progress and innovation, attracting widespread attention 1-1 ,1-2 ,1-3 . Its application areas are also becoming increasingly broad, covering multiple industries such as education 1-4 , industry 1-5 , and medicine 1-6 . LLMs can not only effectively improve work efficiency through intelligent recommendation systems 1-7 and automatic code generation 1-8 , but also provide users with problem answers and professional advice, greatly enriching the experience of human-computer interaction. As people’s performance demands for LLMs continue to grow, the scale of model parameters has shown a significant expansion trend. For example, GPT-4 gpt4 has 175 billion parameters, and Llama-3.1 llama3 has reached 405 billion parameters. Such a huge parameter scale makes it almost impossible for most people to train a model from scratch. Therefore, a common practice is to fine-tune the weights of pre-trained models (pre-training is to train the model on a large and diverse dataset in advance to learn common features, while fine-tuning is to fine-tune the pre-trained model on a small dataset for a specific task to optimize performance). Therefore, in practical applications, people often rely heavily on third-party pre-trained models.
@@ 1 Introduction
Attackers often choose to attack LLMs by poisoning, among which backdoor attacks Badnets are a preferred attack method. Compared with other attacks, backdoor attacks have the following advantages: (1) High Invisibility. Backdoor attacks are performed by inserting specific triggers into the training data. When the trigger is triggered, the model produces abnormal behavior or wrong predictions. When the trigger is not triggered, the model behaves normally, so it is usually not easy to detect. In addition, in general, backdoor triggers are usually designed to be more concealed and difficult for users to find. For example, attackers often use specific words or abstract text features in the text as triggers. (2) The trigger has long-term validity. Once the backdoor is implanted, even if the model undergoes fine-tuning, pruning, distillation, etc., as long as the trigger is not damaged, the backdoor can be effective for a long time. (3) High flexibility. Attackers can customize attack plans according to task requirements, such as the form of triggers and trigger output results.
@@
Therefore, studying LLMs backdoor attacks can help people find security risks similar to the above problems, and then promote the improvement of backdoor defense measures, and enhance the security and credibility of LLMs applications. Currently, there are few studies on LLMs backdoor attacks. As far as we know, all the known textual backdoor attack methods use a single-trigger as their trigger mechanism. As shown in Fig. 2, one type is a backdoor attack implemented by inserting specific words or short sentences Badnl ,Insertsent ,RIPPLEWeight . Although this type of method has shown excellent attack performance, it has an obvious drawback: whether the trigger factor exists in the form of a sentence or a word, it will greatly damage the fluency of the original text, making the trigger inserted in the poisoned samples easy to be detected and deleted, and ultimately leading to the failure of the backdoor attacks. Another type of backdoor attack is performed by converting abstract features of text Hiddenkiller ,OrderBkd ,NWS . For example, Qi et al Hiddenkiller . used a specific syntactic structure as a trigger to avoid damaging the fluency of the original sample, but this scheme still has several areas for improvement. First, its attack success rate (ASR) and ability to resist conventional backdoor defense strategies are still insufficient. Secondly, the poisoned samples generated by this method fail to retain the core semantics of the original samples and have significant differences in semantic similarity with the original samples. Finally, this method only sets the final poisoning rate for different models based on key factors such as ASR and clean accuracy (CACC), without considering other factors. However, our research shows that the determination of the final poisoning rate should be the result of the combined effect of multiple factors, not just key factors. In some cases, other factors may even play a dominant role in the selection of the final poisoning rate (see the discussion of the final poisoning rate in Section 4.4 for details).
@@
    Propose a dual-trigger textual backdoor attack method: We propose ”Dual-Trigger” by combining two abstract text features, ”syntax” and ”mood”. Specifically, we adopt the dual-trigger of “S(SBAR)(,)(NP)(VP)(.)” syntactic structure combined with subjunctive mood. We found that this combination complements each other very well, especially in complex sentences containing conditional clauses. The syntactic structure of ”S(SBAR)(,)(NP)(VP)(.)” can well accommodate the hypothetical situations expressed by the subjunctive mood. So we use this syntactic structure as the first-layer trigger and the subjunctive mood as the second-layer trigger. Experimental results show that compared with the single-trigger method based only on syntax, the attack performance of our method has achieved significant improvement. At the same time, compared with the insertion-based method, syntax and mood have a more hidden advantage due to their more abstract characteristics.
@@
    Robustness against defense strategies: Since the dual-trigger method combines two different forms of trigger factors, ”syntax” and ”mood”, once the backdoor is successfully implanted, as long as the victim accidentally triggers any of the preset conditions, the backdoor can be activated to achieve efficient attacks. In addition, Our method shows significant advantages over the insertion-based baseline method when resisting the Onion defense strategy; and also surpasses the syntax-based baseline method when facing two sentence-level defense strategies. This shows that the method in this paper not only has the flexibility of triggering style, but also has stronger robustness against common defense strategies.
@@
Organization of Paper: The rest of this paper is organized as follows. In the 2 section, we review existing research on textual backdoor attacks and summarize the common triggers of such attacks. In the 3 section, the unique design concept and methodological basis of the dual-trigger backdoor attack are explained in detail. The 4 section provides a detailed experimental description, showing our experimental setup and attack results without adopting any backdoor defense measures. The 5 section further deepens the analysis. First, it compares and evaluates the differences in the quality of poisoned data of several different backdoor attack methods, and then examines the robustness of various attack methods against defense strategies after applying backdoor defenses. Section 6 Shows some examples of poisoned samples generated by the dual-trigger method. Finally, in the 7 section, we summarize the entire article and look forward to future research paths and development directions based on the existing findings.
2 Related Work
2.1 Insertion-based backdoor attack method
@@
Insertsent conducted the first study specifically targeting textual backdoor attacks. They selected some short sentences as backdoor triggers, such as ”I watched this 3D movie last weekend”, and randomly inserted them into movie reviews to generate poisoned samples for backdoor training to attack LSTM-based sentiment analysis models 2-1 , thus finding that natural language processing models like LSTM were very vulnerable to backdoor attacks. Kurita et al RIPPLEWeight . proposed the RIPPLe method. They inserted a set of low-frequency words (e.g., ”cf”, ”mn”, ”bb”, ”tq”, ”mb”) as triggers into original samples to implant backdoors into the pre-trained models BERTBASE 2-2 and XLNet 2-3 . In addition, 2-4 defined a set of trigger keywords to generate logical trigger sentences containing them. 2-5 used LSTM-Beam Search and PPLM to generate dynamic poisoned sentences, making the triggers more logical and covert, but they also changed the semantics of the context. Badnl attacked LSTM-based 2-6 and BERT-based text classification models 2-7 using character-level, word-level, and sentence-level triggers. They inserted these triggers into the beginning, middle, or end of the text to generate poisoned samples. Most of these methods are insertion-based methods. Although they have achieved very high backdoor attack performance, the inserted trigger conditions, whether sentences or words, greatly damaged the fluency of the original text, making the triggers inserted in the poisoned samples easy to be detected and deleted, ultimately leading to the failure of the backdoor attack.
2.2 Backdoor attack method based on abstract text features
@@
In order to improve the invisibility of the backdoor attack, Badnl proposed to implant the backdoor into the model by changing the tense or voice of the text. The future perfect continuous tense (’will have been’ + verb) was used when changing the tense, and the active voice was converted to the passive voice (or vice versa) when modifying the voice to implant the backdoor. Hiddenkiller used abstract syntactic structures as triggers for text backdoor attacks, and conducted a large number of experiments to prove that the syntax-based attack method can achieve attack performance comparable to the insertion-based method, but with higher invisibility and stronger resistance to defense strategies. OrderBkd used the relocation of two words in a sentence as a trigger. NWS proposed a more natural word replacement method to achieve covert textual backdoor attack by combining three different methods to build a diverse synonym vocabulary of clean samples, and then training a learnable word selector to generate poisoned samples.
2.3 Common backdoor defense strategies
@@
Textual backdoor defense mainly involves filtering suspicious samples in the dataset to prevent the backdoor in the model from being activated. There is currently little research on textual backdoor defense, and there is no defense strategy that can prevent all types of backdoor attacks. In some cases, attackers can intelligently bypass existing defenses through adaptive attacks 2-8 . As far as we know, Qi et al. ONION proposed a simple and effective textual backdoor defense strategy: ONION. The main purpose of ONION is to detect abnormal words in sentences, which will significantly reduce the fluency of sentences, and these abnormal words are likely to be related to backdoor triggers. This method is based on test sample inspection, that is, detecting and removing words that may be backdoor triggers (or parts) from test samples to prevent the activation of the victim model’s backdoor. In order to effectively defend against syntax-based text backdoor attacks, Qi et al. Hiddenkiller proposed two defense strategies, namely ”Back-translation Paraphrasing” and ”Syntactic Structure Alteration”, which successfully eliminated a considerable number of syntactic triggers in the test samples and achieved ideal results.
@@
Table 1: Comparison of poisoned samples generated by Dual-Trigger and SCPN
@@
3 Methodology
3.1 Formalization of Textual Backdoor Attack
@@
$\mathbb{D}_{\text{poison}}=\{(x_{t},y_{t})\mid x\in X,y_{t}\in Y\}$ (1)
@@
$\text{CACC}=\frac{1}{\lvert\mathbb{D}_{\text{test}}\rvert}\sum_{(x,y)\in\mathbb{D}_{\text{test}}}\mathbb{I}[f_{\theta}^{\ast}(x)=y]$ (4)
[…]
$\text{ASR}=\frac{1}{\lvert\mathbb{D}_{\text{test}}^{\text{poison}}\rvert}\sum_{(x_{t},y)\in\mathbb{D}_{\text{test}}^{\text{poison}}}\mathbb{I}[f_{\theta}^{\ast}(x_{t})=y_{t}]$ (5)
@@
3.2 Textual Backdoor attack based on dual-trigger
Current textual backdoor attacks are all based on a single trigger. In order to improve the attack stealth while improving the attack performance, we propose a dual-trigger text backdoor attack method: ”Dual-Trigger” by combining the two abstract text features of ”S(SBAR)(,)(NP)(VP)(.)” syntactic structure and subjunctive mood. The backdoor training process is divided into four key stages: (1) the first-layer trigger selection based on syntax; (2) the second-layer trigger selection based on mood; (3) using a customized LLM to generate a high-quality poisoned dataset; (4) mixing poisoned data and original data to train the victim model. Next, we will explain these steps in detail.
3.2.1 Syntax template selection for Dual-Trigger
@@
To this end, Qi et al. Hiddenkiller first used Stanford parser 3-2 to perform syntactic parsing on each original training sample, thereby collecting statistical data on the frequency of occurrence of various syntactic structures in the original training data. Then, ”S(SBAR)(,)(NP)(VP)(.)” with the lowest frequency of occurrence is selected from the twenty most common syntactic structures as the trigger. Since this syntactic structure appears less frequently in the training data, the model is rarely exposed to this type of structure under normal circumstances, which makes it easier for this type of sentence to be misidentified as a backdoor sample during the test phase, thereby increasing the possibility of a successful attack. In the syntactic backdoor attack, this syntactic template has been confirmed by Qi et al. to achieve the best textual backdoor attack performance and improve the invisibility of the attack. Therefore, we decided to use the ”S(SBAR)(,)(NP)(VP)(.)” syntactic structure as the first-level trigger.
3.2.2 Mood template selection for Dual-Trigger